Can you work this problem out for me....In how many ways can a person choose 3 movies to see in a theater playing 11 movies?
Answer by stanbon(75887)

Can you work this problem out for me....In how many ways can a person choose 3 movies to see in a theater playing 11 movies?
-----
1st choice: 11 ways
2nd choice: 10 ways
3rd choice: 9 ways
----
11P3 = 11!/(11-3)! = 11*10*9 = 990
=====================================
Cheers,
Stan H.